OpenStudy (anonymous):

What are the positive residual obtained when (17 ^ 15 +8 ^ 25) ^ 10 divided by 10?

OpenStudy (anonymous):

should be 1

OpenStudy (anonymous):

do you want an explanation?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

ok :) so, to start off, we need to understand that the positive residual is like a remainder...so anything that is a remainder after being divided by 10 is just the number that is in the one's digit....

OpenStudy (anonymous):

like....225 divided by 10...the residue would be 5. 332 divided by 10, the residual would be 2. Do you get that??

OpenStudy (anonymous):

Maribor?

OpenStudy (anonymous):

can you show me how you resolved the issue

OpenStudy (anonymous):

I will, but I need to make sure you got what I said above XD

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

ok, good :) so that means you only need to concentrate on the last digit of every number. 17^1=17 last digit is 7 17^2=289 last digit is 9 17^3=4913 last digit is 3 17^4=83521 last digit is 1 17^5=1419857 last digit is 7 So, by this, we can see that this sequence of 7931 is going to loop every four exponents. In the question above, 17 is raised to the 15 th power. Therefore, 7 9 3 1 7 9 3 1 7 9 3 1 7 9 3 The 15th number ends with 3.

OpenStudy (anonymous):

OK Thanks for your help

OpenStudy (anonymous):

Now, taking the same pattern with 8: 8 4 2 6 8 4 2 6 8 4 2 6 8 4 2 6 8 4 2 6 8 4 2 6 8. The last number raised to the 25th power is 8.

OpenStudy (anonymous):

add these last digits up and you should get 1. 1^10 is 1 so the remainder of this equation is 1. Sorry for the lengthy explanation XD

OpenStudy (anonymous):

Thanks again